# V2X Cooperative Perception for Autonomous Driving: Recent Advances and Challenges

### v3

@@ IV-D3 Attention-based Fusion
To address the above challenges, SCOPE [194] introduced a Context-aware Information Aggregation (CIA) approach to incorporate spatiotemporal information from previous frames of the ego agent. A pivotal component of CIA is a pyramid LSTM, which replaces the conventional matrix multiplication of LSTM with a multi-scale convolutional structure. The resultant hidden state from the pyramid LSTM serves as a refined context-aware feature. This feature and collaboration and ego features are fused based on their complementary contributions and spatial attention maps obtained through max pooling. However, it is worth noting that max-pooling may not fully capture the underlying correlations among informative features. In contrast, IoSP-CP [156] devised a Historical Prior Hybrid Attention (HPHA) fusion algorithm that leverages a multi-scale transformer and a short-term attention module. HPHA assigns varying weights to different information from collaborators to optimize feature fusion. The multi-scale transformer module extracts attention weights at different scales, facilitating collaborative correlation learning across distinct spatial dimensions. The short-term attention mechanism captures temporal dependencies and identifies discriminative object features from the temporal dimension. HPHA empowers IoSP-CP to enhance informative features while reducing background noise.
@@
TABLE VII: Summary of the Works on Issues of Cooperative Perception and Corresponding Methods.
@@
MACP [222] 2023 V2V LiDAR V2V4Real, OPV2V Intermediate Detection $\surd$
@@
DI-V2X [224] 2023 V2X LiDAR DAIR-V2X, V2XSet Intermediate Detection $\surd$
@@
DRR MWV2IComm[227] 2021 V2I -- -- -- Passing intersection $\times$
@@
V2X-INCOP [229] 2023 V2X LiDAR V2X-Sim,OPV2V,DAIR-V2X Intermediate Detection $\times$
DT TrustEsti [230] 2019 V2V Camera/LiDAR -- -- -- $\times$
@@
CAD [232] 2023 V2V LiDAR Adv-OPV2V, Adv-MCity Early/Intermediate Detection $\surd$
MADE [233] 2023 V2V LiDAR V2X-sim, DAIR-V2X Intermediate Detection $\times$
@@
MOT-CUP [237] 2023 V2X LiDAR V2X-Sim Late Detection, Tracking $\times$
DMSTrack [238] 2024 V2V LiDAR V2V4Real Late Tracking $\times$
@@
FedBEVT [242] 2023 V2X Camera CARLA&OpenCDA -- Detection $\surd$
FL-CCS [243] 2023 V2X -- -- -- -- $\times$
S2R S2R-ViT [244] 2023 V2V LiDAR OPV2V,V2V4Real Intermediate Detection $\times$
@@
  • Note: MH: Model Heterogeneity, DH: Data Heterogeneity, DRR: Data Rate Requirements, LC: Lossy Communication, DT: Data Trustworthiness, PU: Perception Uncertainty, TD: Task Discrepancy, DP: Data Privacy, S2R: Simulation to Reality.
IV-D4 Hybrid Fusion
Different collaboration schemes have advantages and limitations and can complement each other effectively. However, single-type fusion methods may not fully exploit the potential for deep fusion of various data types and generating higher-level collaborative representations. Hybrid fusion-based approaches offer promise in harnessing the strengths of diverse collaboration schemes by integrating multiple fusion methods within a single CP model.
Typically, objects near the onboard sensor exhibit a high point density and are more likely to be detected using data from a single sensor. Conversely, distant points are often sparse due to limited sensor visibility. Therefore, employing an early fusion scheme in areas far from the Ego can enhance object detection likelihood. This concept has been validated by research [110], which introduced a V2I cooperative 3D object detection model using a hybrid fusion approach to aggregate point cloud data from multiple infrastructure sensors. The central idea behind this hybrid fusion strategy is to transmit raw data (early fusion) in regions distant from a sensor where point cloud density is low and to share object-level information (late fusion) in the immediate vicinity of a sensor. This approach proves cost-effective for autonomous driving systems in complex road segments but may lack flexibility in adjusting the proportion of raw and object-level data.
In contrast, ML-Cooper [216] employed an improved Soft Actor-critic (K-SAC) [246] algorithm to dynamically reduce channel congestion. It combines three distinct cooperation schemes to maximize V2V network bandwidth utilization. Specifically, each CAV partitions its sensing data into three segments based on point cloud angle and density and shares raw, feature, and object data from each segment with other CAVs. The K-SAC algorithm determines the distribution among the three segments based on the prevailing channel conditions. This adaptability characteristic equips ML-Cooper to operate effectively in highly dynamic vehicular networks.
@@ V Issues of Cooperative Perception and Corresponding Methods
While the majority of research on CP for driving automation has focused on key aspects of the collaboration process, such as communication latency, information alignment, and fusion strategies, as discussed in the previous section, limited attention has been given to subtle yet crucial challenges that have practical implications for advancing Cooperative ITS. This section delves into exploring recent findings and the corresponding methodologies to address these practical challenges. Table VII provides a comprehensive overview of characteristics related to these advancements.
@@ V-A Agent Heterogeneity
Although CP systems have made significant progress, it is important to acknowledge that these improvements often assume that the agents involved are uniform. In reality, it is difficult to achieve uniformity in agent configuration and sensor equipment across connected agents, including CAVs and infrastructures, due to varying preferences among system developers and automobile manufacturers. Therefore, it is necessary to prioritize finding solutions to the practical challenges of enabling effective collaboration among diverse agents within V2X systems.
@@ V-A1 Model Heterogeneity
Object-level-based cooperative detection frameworks can circumvent the necessity for comprehensive model data exchange by exchanging perception outcomes, including bounding boxes and confidence scores, among participating agents [219]. Nevertheless, due to model heterogeneity, the confidence scores provided by different agents may not always align accurately. In [219], a model-agnostic CP framework was proposed to tackle this confidence misalignment problem. This framework employed a customized confidence calibrator called Doubly Bounded Scaling (DBS) and a tailored Promote-suppress Aggregation (PSA) algorithm. DBS learned a scaling function based on the Kumaraswamy Cumulative Density Function [247] to eliminate confidence score bias. At the same time, PSA selected high-score bounding boxes based on calibrated confidence and spatial consistency of the same object for bounding box fusion. This approach underscored the importance of output calibration across agents with different perception models for effective late collaboration.
In addition to variations in confidence levels, connected agents’ detection outcomes may contain different information due to differences in their motion prediction models. For instance, some agents may not provide information about detected objects’ acceleration and yaw rate, resulting in heterogeneous collaborative perception fusion with varying state spaces. In [220], researchers introduced a CI-based track-to-track fusion approach to address tracks with unequal dimensions. This approach involved augmenting the low-dimensional track and exploring various methodologies for state augmentation. The effectiveness of state augmentation significantly impacted the CI’s optimization process, as demonstrated through comparative analysis.
Late collaboration can maintain collaborators’ model information privacy by allowing each agent to train its model individually and share outputs. However, individual outputs may introduce noise, leading to a degradation in cooperative performance. Intermediate collaboration has become a prominent choice for V2X CP due to its flexibility. In real-world scenarios, connected agents may use different feature extractors, resulting in a domain gap among shared features. To mitigate this feature gap, the work in [221] proposed a Multi-agent Perception Domain Adaptation (MPDA) framework. The MPDA framework included a feature resizer and a sparse cross-domain transformer. The feature resizer aligned the spatial and channel dimensions of received intermediate features using the FAX attention block [215] and a convolution layer. The sparse cross-domain transformer generated domain-invariant features through an adversarial strategy to deceive a domain classifier, effectively reducing the feature domain disparities when connected agents employed different feature extraction networks.
Drawing inspiration from the parameter-efficient fine-tuning strategy [248], a Model Adaptation for Cooperative Perception (MACP) was proposed in [222]. Specifically, a single-agent pre-trained model was adapted to the cooperative setting by freezing most of its parameters and adding lightweight modules. Domain shift mitigating was achieved through the design of the Convolution Adapter (ConAda) for the feature encoder and communication channel. The ConAda module helps adapt the pre-trained single-agent parameters to the cooperative setting by projecting the source feature map to a lower dimension and then transforming it back to the source space. This adaptation allows the model to handle spatial distribution shifts and feature space shifts that occur when combining point clouds from different vehicles. The MACP framework effectively addressed domain shifts and enhanced the perception capabilities. However, it may suffer from biases in the pre-trained model, leading to lower average precision for objects in the 50-100m range compared to other methods.
@@ V-A2 Data Heterogeneity
In practical autonomous driving scenarios, connected agents often possess distinct sensory modalities. Agents with LiDAR sensors provide precise and accurate geometrical information, while those equipped with cameras offer comprehensive and contextually rich semantics. Collaborating with agents possessing diverse sensor modalities can enhance the ego’s sensing capabilities by leveraging unique yet complementary environmental attributes captured by each sensor. However, sharing information among camera and LiDAR agents introduces semantic disparities. Additionally, disparities exist between measurements from roadside infrastructures and those from CAVs. Different sensors may exhibit varying noise levels due to factors like maintenance frequency and hardware precision. These distinctive attributes of connected agents present significant challenges in developing a heterogeneous cooperative system.
@@
In a multi-agent hetero-modal setting, connected agents are equipped with various sensors, forming a dynamically heterogeneous graph. The presence and types of these sensors are stochastic, and their relative poses change across different scenes. This necessitates an adaptive architecture that can accommodate diverse agent quantities and typologies while addressing the misalignment of CP data. To address these challenges, HM-ViT [223] facilitated V2V cooperative 3D detection through a Heterogeneous 3D Graph Attention (H3GAT) module. In this module, each node represents a feature vector of an agent’s feature map, and its core function is to extract heterogeneous inter-agent and intra-agent visual cues within the BEV feature space. Node states are updated by aggregating features derived from transformer blocks with a heterogeneous normalization layer, followed by a hetero-modal Multi-layer Perceptron Layer that processes camera and LiDAR features with distinct parameter sets. HM-ViT demonstrated remarkable flexibility and robustness while achieving superior performance in scenarios with varying sensor modalities.
Sensors of identical modalities may exhibit heterogeneity in terms of brand, resolution, and configuration parameters. The overlook of the domain discrepancy engendered by the disparate sensor attributes utilized by each entity can deleteriously impact the culminating efficacy of the system [224]. Consequently, it becomes imperative to devise a mechanism for the extraction of domain-agnostic representations from data originating from multiple sources within the ambit of V2X collaborative perception.
In response to this exigency, a teacher-student distillation architecture, designated as DI-V2X, was introduced in [224]. This framework is characterized by the integration of three principal components: a Domain-Mixing Instance Augmentation (DMA) module, a Domain-Adaptive Fusion (DAF) module, and a Progressive Domain-Invariant Distillation (PDD) module. The DMA module was deployed to mediate the distributional alignment between the teacher model and the student models, thereby laying the groundwork for an efficacious distillation endeavour. Concurrently, the DAF module amalgamates features extracted from both vehicles and infrastructure, engendering a composite representation enriched with pertinent information from each respective domain. In a parallel vein, the PDD module orchestrates the alignment of disparate student features with the teacher’s BEV feature, both before and after fusion, under the directive of overlapping and non-overlapping masks. Through the synergetic operation of these modules, DI-V2X substantiated an enhancement in V2X perception tasks, fostering a generalized representation that reconciles point cloud data across a spectrum of domains.
@@
Besides sensor characteristics, diverse environmental conditions inevitably create a substantial domain gap and data heterogeneity among CAVs [225]. To address these issues, the work in [225] proposed a domain generalization framework comprising an Amplitude Augmentation (AmpAug) method, an intra-system domain alignment mechanism, and a meta-consistency training scheme. The AmpAug is to increase the diversity of low-frequency variations, forcing the model to obtain domain-invariant representations. This enhancement can reduce the model’s influence on low-spectral variations in the image. The meta-consistency training scheme was utilized to simulate domain shifts and optimize the model with consistency loss. The domain alignment mechanism translates image styles and unifies the distribution of image pixels in colour space among CAVs before the inference stage, potentially eliminating the domain discrepancy among CAVs prior to inference.
Furthermore, the distinctness in data characteristics caused by different types of agents also poses significant challenges in designing a heterogeneous cooperative system. VINet [226] addressed this issue using a Two-stream Fusion (TSF) algorithm to integrate heterogeneous features captured from infrastructures and CAVs on an edge server. Specifically, a Lightweight Feature Extraction (LFE) module reduced computing loads in feature extraction without involving the CNN backbone. TSF employed two feature streams with different LFEs to extract pillar features from CAVs and connected infrastructures. Each feature stream implemented a feature mapping and max-out procedure to calculate fused features. The resulting fused features from the two streams were concatenated and interpreted by a CNN backbone to generate detected bounding boxes. VINet proved effective for cooperative object detection with low system-wide costs under dynamic, scalable, and heterogeneous conditions.
V-B Data Rate Requirements
The data rate requirements for CP are expected to increase significantly with the advancement of automated driving levels [249]. Reliable communication is a critical necessity for CDA, as it involves continuously and timely updating dynamic information about surrounding moving obstacles to prevent traffic accidents. Therefore, determining the minimum required sensor data rate for ensuring safe autonomous driving becomes imperative.
Organizations like the 3rd Generation Partnership Project (3GPP) and the 5G Automotive Association (5GAA) have presented data rate requirements to support V2X CP-related applications [250, 251]. However, these considerations did not account for safety concerns such as cyber-attacks and software errors. In a study conducted in [227], the minimum required data rate for a CAV to safely navigate through an intersection was analyzed in scenarios involving both automated and human driving. Safety was assessed by distinguishing cases where braking is necessary to prevent a collision, thereby reducing unnecessary braking while ensuring safe driving and traffic efficiency. The minimum required data rate for each velocity category was determined based on case classification and compared with the data rate achievable through conventional V2I and millimeter-wave communication bands. The findings indicated that transmission using the new millimetre-wave communication standards could meet the minimum data rate requirement for safe autonomous driving in such scenarios.
V-C Lossy Communication
In real-world traffic scenarios, lossy communication is common in V2X communication due to various factors. These factors include Doppler shift induced by fast-moving vehicles [252], dynamic topology leading to routing failures [253], adverse weather conditions, and obstacles causing multi-path effects and signal power fluctuations, resulting in packet loss [254]. Additionally, shared data may face interference from other communication networks or be tampered with by attackers before reaching their intended destination. These factors collectively contribute to incomplete or inaccurate shared information during V2X communication, consequently diminishing the performance of CP systems.

### v4

@@ IV-E3 Attention-based Fusion
The aforementioned attention-based models have significantly advanced cooperative information fusion by focusing primarily on spatial dimensions. However, their reliance on single-frame prediction neglects historical context and overlooks important temporal semantic cues, limiting their ability to detect fast-moving objects due to point cloud sparsity. To address these limitations, SCOPE [196] introduced a Context-aware Information Aggregation model that uses a pyramid LSTM to incorporate spatiotemporal data from previous ego agent frames, generating refined context-aware features. These features are fused with cooperative features based on complementary contributions and spatial attention maps obtained through max pooling, which may not fully capture correlations among informative features. To overcome this, IoSP-CP [121] proposed the Historical Prior Hybrid Attention (HPHA) fusion algorithm, comprising a multi-scale transformer and a short-term attention module. The multi-scale transformer extracts attention weights at different spatial scales, while the short-term attention mechanism captures temporal dependencies and identifies discriminative features, enhancing informative features while reducing background noise.
@@
TABLE VI: Summary of the Works on Issues of Cooperative Perception and Corresponding Methods.
@@
MACP [228] 2024 V2V LiDAR V2V4Real, OPV2V Intermediate Detection $\surd$
@@
DI-V2X [229] 2024 V2X LiDAR DAIR-V2X, V2XSet Intermediate Detection $\surd$
@@
V2X-M2C [232] 2024 V2X LiDAR V2XSet, OPV2V Intermediate Detection $\times$
VIMI [233] 2023 V2I Camera DAIR-V2X Intermediate Detection $\surd$
MH & DH HEAL [234] 2024 V2X Camera&LiDAR OPV2V-H, DAIR-V2X Intermediate Detection $\surd$
@@
V2X-INCOP [236] 2024 V2X LiDAR V2X-Sim,OPV2V,DAIR-V2X Intermediate Detection $\times$
SsAW [237] 2024 V2V LiDAR OPV2V, V2V4Real Intermediate Detection $\times$
TempCoBEV [238] 2024 V2V Camera OPV2V Intermediate Segmentation $\times$
CD CooperFuse [239] 2024 V2X LiDAR V2X-real Late Detection $\times$
FFNet [240] 2023 V2I LiDAR DAIR-V2X Intermediate Detection $\surd$
SyncNet [241] 2022 V2V LiDAR V2V-Sim Intermediate Detection $\surd$
V2VNet [140] 2020 V2V LiDAR V2V-Sim Intermediate Detection, Motion Prediction $\times$
V2X-ViT [145] 2022 V2X LiDAR V2XSet Intermediate Detection $\surd$
V2X-PC [136] 2024 V2X LiDAR DAIR-V2X, V2XSet Early Detection $\times$
FreeAlign [195] 2024 V2X LiDAR OPV2V, DAIR-V2X Intermediate Detection $\surd$
LC & CD Ro-temd [243] 2024 V2V LiDAR OPV2V Intermediate Detection $\surd$
DS AdversAttack [244] 2021 V2V LiDAR V2V-Sim Intermediate Detection $\times$
ROBOSAC [245] 2023 V2X LiDAR V2X-Sim Intermediate Detection $\surd$
TrustEsti [246] 2019 V2V Camera/LiDAR -- -- -- $\times$
@@
CAD [248] 2024 V2V LiDAR Adv-OPV2V, Adv-MCity Early/Intermediate Detection $\surd$
MADE [249] 2024 V2V LiDAR V2X-sim, DAIR-V2X Intermediate Detection $\times$
FedBEVT [251] 2023 V2X Camera CARLA&OpenCDA -- Segmentation $\surd$
FL-CCS [252] 2023 V2X -- -- -- -- $\times$
FedDWA [253] 2024 V2X Camera CARLA&OpenCDA -- Segmentation $\times$
@@
MOT-CUP [258] 2024 V2X LiDAR V2X-Sim Late Detection, Tracking $\times$
DMSTrack [259] 2024 V2V LiDAR V2V4Real Late Tracking $\surd$
GCDA [260] 2024 V2V LiDAR V2V4Real Late Tracking $\times$
@@
S2R S2R-ViT [263] 2024 V2V LiDAR OPV2V,V2V4Real Intermediate Detection $\times$
@@
  • Note: MH: Model Heterogeneity, DH: Data Heterogeneity, LC: Lossy Communication, CD: Communication Delays, DS: Data Security, DP: Data Privacy, PU: Perception Uncertainty, TD: Task Discrepancy, S2R: Simulation to Reality.
IV-E4 Hybrid Fusion
Different collaboration schemes have advantages and limitations and can complement each other effectively. However, single-type fusion methods may not fully exploit the potential for both deep fusion of various data types and consuming minimum communication bandwidth by generating higher-level collaborative representations. Hybrid fusion-based approaches offer promise in harnessing the strengths of diverse collaboration schemes by integrating multiple fusion methods within a single CP model, thus balancing the accuracy and transmission data size.
Objects near the onboard sensor typically exhibit a high point density and are more easily detected using data from a single sensor, while distant points tend to be sparse due to limited sensor visibility. Early collaboration for distant regions can improve detection accuracy, as demonstrated by research [220], which proposed a V2I cooperative 3D object detection model using hybrid fusion. This method transmits raw data (early fusion) for low-density areas and shares object-level information (late fusion) for nearby areas, providing a cost-effective solution but with limited flexibility in adjusting data sharing. Hybrid CP [222] proposed a region-based hybrid collaborative perception strategy. Collaboration is classified into two types based on the overlap between the detection ranges of vehicles. For the overlapping area, feature-level intermediate collaboration is used by exchanging and fusing features from various CAVs. Late collaboration is carried out in the non-overlapping area by developing and sharing the local detection result with a low communication volume. Conversely, ML-Cooper [221] employs an improved Soft Actor-Critic [265] algorithm to optimize V2V bandwidth by dynamically adjusting data sharing among raw, feature, and object levels based on channel conditions, enhancing adaptability in dynamic vehicular networks.
@@ V Issues of Cooperative Perception and Corresponding Methods
While the majority of research on CP for driving automation has focused on key aspects of the collaboration process, such as communication latency, information alignment, and fusion strategies, as discussed in the previous section, limited attention has been given to subtle yet crucial challenges that have practical implications for advancing Cooperative ITS. This section delves into exploring recent findings and the corresponding methodologies to address these practical challenges. Table VI provides a comprehensive overview of characteristics related to these advancements.
@@ V-A Agent Heterogeneity
Although CP systems have made significant progress, it is important to acknowledge that these improvements often assume that the agents involved are uniform. In reality, it is difficult to achieve uniformity in agent configuration and sensor equipment across connected agents due to varying preferences among system developers and automobile manufacturers. Therefore, it is necessary to prioritize finding solutions to the practical challenges of agent heterogeneity, including model and data discrepancies among agents.
@@ V-A1 Model Heterogeneity
Object-level CP frameworks avoid the need for full model data exchange by sharing perception results like bounding boxes and confidence scores participating agents [225]. Nevertheless, due to model heterogeneity, the confidence scores provided by different agents may not always align accurately. In [225], a model-agnostic CP framework was proposed to tackle this confidence misalignment problem, using a Doubly Bounded Scaling calibrator to eliminate bias in confidence scores and a Promote-suppress Aggregation algorithm to select high-score bounding boxes based on calibrated confidence and spatial consistency. In addition to variations in confidence levels, connected agents’ detection outcomes may vary due to differences in motion prediction models, with some agents lacking information like acceleration or yaw rate. To address heterogeneous perception fusion with varying state spaces, a CI-based track-to-track fusion approach was proposed in [226], augmenting low-dimensional tracks and exploring methods for state augmentation.
Late collaboration preserves model privacy by allowing agents to train models individually and share outputs, but this can introduce noise and degrade performance. Intermediate collaboration is more flexible and popular for V2X CP, but feature extractors’ differences can create a domain gap. To overcome this issue, a Multi-agent Perception Domain Adaptation (MPDA) framework was introduced in [227], incorporating a feature resizer and a sparse cross-domain transformer to align and produce domain-invariant features, thereby minimizing disparities between agents utilizing heterogeneous feature extraction networks. Inspired by the parameter-efficient fine-tuning strategy [266], MACP [228] introduced a Convolution Adapter (ConAda) to mitigate domain shifts by projecting the source feature map to a lower dimension and then transforming it back. Additionally, a single-agent pre-trained model was adapted to cooperative settings by freezing most parameters and adding lightweight modules. While this approach effectively mitigates domain shifts and enhances perception, it may exhibit reduced precision for objects in the 50-100m range due to inherent biases in the pre-trained model.
@@ V-A2 Data Heterogeneity
In practical autonomous driving, connected agents often have different sensor modalities: LiDAR provides precise geometry, while cameras offer rich semantics. Collaborating with agents with diverse sensors enhances perception by leveraging complementary information. However, sharing data between camera and LiDAR agents introduces semantic disparities, and measurements differ between roadside infrastructure and CAVs. Additionally, sensor noise varies due to maintenance and hardware precision, posing challenges for developing a heterogeneous cooperative system.
@@
In a multi-agent hetero-modal setting, connected agents have various sensors, forming a dynamic heterogeneous graph with stochastic sensor types and changing poses across scenes. To handle these challenges, HM-ViT [164] introduced the H3GAT module for V2V cooperative 3D detection, which updates node states by extracting inter-agent and intra-agent cues using transformer blocks and a hetero-modal MLP layer. This allows processing camera and LiDAR features with distinct parameters, enabling flexibility, robustness, and superior performance in scenarios with diverse sensor modalities.
On the other hand, identical modality sensors can vary in brand, resolution, and configuration, leading to domain discrepancies that impact system performance. In response to this exigency, DI-V2X [229] introduced a teacher-student distillation architecture with three key components: the Domain-Mixing Instance Augmentation module for aligning distributions between teacher and student models, the Domain-Adaptive Fusion module for combining vehicle and infrastructure features, and the Progressive Domain-Invariant Distillation module to align student and teacher features before and after fusion. This method improves V2X perception by creating a generalized representation across various domains.
@@
Besides sensor characteristics, diverse environmental conditions inevitably create a substantial domain gap and data heterogeneity among CAVs [230]. To address this, [230] proposed a domain generalization framework with an amplitude augmentation module to reduce the model’s influence on low-spectral variations, a domain alignment mechanism to unify image pixel distributions, and a meta-consistency training scheme to simulate domain shifts and optimize the model with consistency loss, potentially eliminating the domain discrepancy among CAVs prior to inference.
The distinctness in data characteristics across different agent types presents significant challenges in CP. VINet [231] addressed this challenge using a Two-stream Fusion algorithm to integrate heterogeneous features from infrastructures and CAVs via a feature mapping and max-out procedure on an edge server. In V2X-M2C [232], a heterogeneity-reflected convolution module was introduced to manage multiple heterogeneous agents using two parallel convolution pipelines, considering their different feature distributions while maintaining a lightweight and efficient architecture. However, time asynchrony and calibration errors between vehicle and infrastructure cameras can lead to inaccurate relative position detection, causing misalignment between ground truth and projected 2D bounding boxes. VIMI [233] tackled the calibration noise issue through its Multi-scale Cross Attention and Camera-aware Channel Masking modules, using multi-scale feature fusion and camera parameter priors to correct for misalignment caused by time asynchrony.
V-A3 Model and Data Heterogeneity
While some existing studies address the heterogeneity of data and models, effectively modelling dynamically increasing heterogeneous scenarios continues to pose significant challenges. In reality, new heterogeneous agent types with different modalities and models may continuously emerge, causing domain gaps and limiting scalability. HEAL [234] solved this problem by establishing a unified feature space for the initial homogeneous agents and then aligning new heterogeneous agents to this space through a novel reverse alignment mechanism. This approach enables the integration of new agent types with minimal training costs while preserving model and data privacy. Despite its scalability and efficiency, HEAL’s reliance on BEV features requires compatibility with keypoint-based LiDAR models, and its performance depends on accurate agent localization, which can be affected by localization noise in real-world scenarios.
V-B Lossy Communication
In practical V2X communication, factors such as interference, tampering, Doppler shifts [267], routing failures [268], multipath effects, and signal fluctuations [269] frequently cause incomplete or inaccurate information exchange, thereby diminishing the performance of CP.